# Optimal Control of Reserve Asset Portfolios for Pegged Digital Currencies

### v1

@@ 4.3 Presentation Of Results
Revenue is calculated as the negative of the cost functional. Table 1 shows the total revenue by scenario and control strategy. Table 2 displays the fallout rates and responsiveness measures. We simulated 100 scenarios and report average values for all metrics, except for the fallout rate, which is shown as a relative frequency. Figure 1 illustrates the trajectories of the control input $\omega$ for all analyzed control strategies and scenarios.
Strategy Scenario Total Revenue
Optimal Control Single Shock $1.86\text{\times}{10}^{8}$
Prolonged Clustering $1.20\text{\times}{10}^{8}$
False Alarm $1.22\text{\times}{10}^{8}$
Maximum Yield Single Shock $5.52\text{\times}{10}^{7}$
Prolonged Clustering $3.01\text{\times}{10}^{7}$
False Alarm $7.85\text{\times}{10}^{7}$
@@
Table 1: Total Cost by Strategy and Scenario
Strategy Scenario Frequency Resp. (days)
Optimal Control Single Shock $0.01$ 1.75
Prolonged Clustering $0.01$ 2.47
@@
Maximum Yield Single Shock 11 0
Prolonged Clustering 11 0
False Alarm 11 n/a
@@
(a) Single Shock
(b) Prolonged Clustering
(c) False Alarm
Figure 1: Control actions under different scenarios. The y-axis is scaled in millions dollar per hour. The optimal control is color-coded in green, the maximum-yield strategy in gray, and the maximum-liquidity strategy in blue. The maximum-liquidity strategy initially shifts massively into T-bonds but must later sell these positions to meet redemption requests. The optimal control strategy behaves cautiously at first and increases its T-bond investments only after the stress period. The maximum-yield strategy, by definition, makes no reallocations.
The simulation results demonstrate the significant advantages of the proposed control strategy over both myopic yield maximization (maximum yield strategy) and the conservative cash-only strategy. The maximum yield strategy proves to be clearly inferior: this approach leads to excessive initial bond purchases across all simulated scenarios, and systematically confronts the issuer with high redemption obligations and associated capitalization problems toward the end of the simulation period. Consequently, this strategy exhibits clear underperformance compared to the optimized control both in terms of generated revenue and default rates. The responsiveness of this strategy exceeds that of the the proposed approach, as the maximum yield controller usually already experiences liquidity constraints at the time of shock onset and consequently must react more promptly to increased redemption requests. The advantages of our method over the maximum liquidity strategy are more moderate. However, particularly during the more stable market phase at the end of the simulation period, it becomes evident that larger-volume bond purchases are justifiable. Through this strategic orientation, the optimized controller achieves a significant financial advantage over the conservative strategy.
@@ 5 Conclusion
This paper develops an operations-first framework for managing the reserves of a pegged digital currency, accounting for clustering dynamics in order and redemption flows and an endogenous dynamical relationship between peg deviation and redemption intensities. We represent mints and redemptions as self-exciting point processes and treat peg behavior as an explicit operational outcome by directly linking price deviations to cash shortfalls within specified time windows. The coin issuer’s control mechanisms comprise two key instruments: the reallocation rate between immediately available cash and short-duration government bills, and the spread applied to mint or burn requests.
@@
The numerical experiments show how these design choices translate into outcomes that matter for issuers. Relative to a maximum-yield benchmark that waits until the last moment to raise liquidity, and a maximum-liquidity benchmark that forgoes carry, the window-based controller preserves most of the bill yield in calm conditions while moving earlier and more decisively when redemption risk rises. Because the policy depends only on filtered intensities, outstanding supply, and window-averaged shadow values, it fits daily treasury workflows and supports supervisory review. More broadly, the framework connects balance-sheet choices to market observables in a way that is transparent: peg deviations emerge only when cash on hand is insufficient within a window, and the controller’s reallocation speed and fee choices directly manage that risk.
A critical risk management concern is the tail probability of a complete depeg. Since such an event is close to catastrophic for the issuer, its probability warrants careful measurement even when it is very small. In this article we simulated 100 runs, which provides only limited estimation precision of probabilities below the one percent level. Future analyses can tighten this by running much larger simulation batches on specialized hardware, by applying dedicated rare-event techniques, or by developing analytical tail approximations.
@@
The authors would like to thank TUM Blockchain Club for insightful discussions that helped shape the content of this manuscript.
@@
The authors report there are no competing interests to declare.

### v2

@@ 4.3 Presentation Of Results
Revenue is calculated as the negative of the cost functional. Table 1 shows the total revenue by scenario and control strategy. Table 2 displays the fallout rates and responsiveness measures. We simulated 100 scenarios and report average values for all metrics, except for the fallout rate, which is shown as a relative frequency. Figures 13 display the peg deviations, and Figures 46 illustrate the trajectories of the control input $\omega$ for all analyzed control strategies and scenarios.
Strategy Scenario Total Revenue [\$M]
Optimal Control Single Shock 98.02
Prolonged Clustering 421.34
False Alarm 300.79
Maximum Yield Single Shock 83.17
Prolonged Clustering 100.61
False Alarm 296.41
@@
Equal Weights Single Shock 51.03
Prolonged Clustering 213.33
False Alarm 151.03
Table 1: Avg. Revenue by Strategy and Scenario
Strategy Scenario Depeg Frequency Resp. (days)
Optimal Control Single Shock 0 2.77
Prolonged Clustering 0 0.67
@@
Maximum Yield Single Shock 1 0.33
Prolonged Clustering 1 0.33
False Alarm 0 n/a
@@
Equal Weights Single Shock 0.69 0.33
Prolonged Clustering 0 6.00
False Alarm 0 n/a
@@
Figure 1: Peg deviation $\Delta P(t)$ under the single-shock scenario for all four control strategies. The shaded region marks the stress window (days 30-45).
Figure 2: Peg deviation $\Delta P(t)$ under the prolonged-clustering scenario for all four control strategies. The shaded region marks the stress window (days 30–70).
Figure 3: Peg deviation $\Delta P(t)$ under the false-alarm scenario for all four control strategies. The shaded region marks the stress window (days 30–35).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Reallocation rate $\omega(t)$ under the single-shock scenario for all four strategies.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Reallocation rate $\omega(t)$ under the prolonged-clustering scenario for all four strategies.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Reallocation rate $\omega(t)$ under the false-alarm scenario for all four strategies.
The simulation results demonstrate the effectiveness of the proposed control strategy across all three stress scenarios. The optimal controller adapts to changing market regimes without prior knowledge of when stress transitions occur, relying solely on filtered intensity estimates and forward-looking costate evaluation. This is visible in Figures 46: the reallocation rate shifts decisively toward cash building as redemption pressure materializes, and reverts to yield-seeking allocations once intensities subside. By contrast, the three benchmark strategies each exhibit characteristic failure modes that illustrate the limitations of static allocation rules.
The maximum yield strategy proves clearly inferior across all scenarios. Its control trajectory (Figures 46, upper right panels) reveals persistent high-frequency oscillation between maximum cash-to-bill and bill-to-cash transfers, driven by the strategy’s reactive logic of liquidating bills only when cash is immediately insufficient. This oscillation is operationally costly and prevents sustained accumulation of yield-bearing reserves. More critically, aggressive yield building at stress onset delays the cash build-up needed to absorb clustered redemptions: once the accumulated liquidity shortfall triggers self-exciting bank-run dynamics via the peg-redemption feedback loop (equation (17)), the resulting depeg is severe and irreversible within the simulation horizon (Figures 12). The economic consequences are twofold: direct peg-deviation costs and a permanent reduction of total investable reserves through realized redemptions at distressed levels. In the prolonged-clustering scenario, this failure emerges later (around day 50) as the near-critical branching ratio sustains aftershock bursts that eventually overwhelm the depleted cash buffer.
The maximum liquidity strategy eliminates depeg risk entirely but forgoes all bill carry, yielding zero revenue across all scenarios. The equal weights strategy offers a partial improvement over maximum yield, maintaining a 50/50 target that provides a larger initial cash buffer. However, its backward-looking rebalancing rule does not adapt the target allocation to prevailing or anticipated stress conditions. In the single-shock scenario, the controller operates near maximum cash-building rate throughout the stress window, yet the 50/50 target tracks a shrinking reserve base as net redemptions accumulate, so the absolute cash level targeted also declines. The reallocation rate does not exhibit a pronounced downward shift at the depeg onset (Figure 4, lower right) precisely because the controller is already at or near its maximum feasible cash-building speed—there is no additional capacity to deploy. This proves insufficient under sustained pressure, resulting in severe depegs in 69% of single-shock runs. In the prolonged-clustering and false-alarm scenarios, the equal weights strategy avoids depegs, but its fixed allocation rule prevents it from capturing the higher bill carry that the optimal controller achieves during calm phases and during stress episodes where full cash conversion is unnecessary. The optimal controller is thus the only strategy that simultaneously (i) builds yield-bearing reserves efficiently during calm periods, as evidenced by positive reallocation rates in the pre-stress and post-stress phases of Figures 4–6, and (ii) achieves sufficiently swift and deep reallocation to liquid cash under stress to prevent the self-reinforcing depeg dynamics that afflict both the maximum yield and equal weights strategies. The false-alarm scenario (Figure 3) further confirms that this responsiveness does not come at the cost of overreaction: the controller absorbs the transient mint surge and modest redemption increase without unnecessary cash accumulation, preserving carry throughout.
@@ 5 Conclusion
This paper develops an operations-first real-time framework for managing the reserves of a pegged digital currency, accounting for clustering dynamics in order and redemption flows and an endogenous dynamical relationship between peg deviation and redemption intensities. We represent mints and redemptions as self-exciting point processes and treat peg behavior as an explicit operational outcome by directly linking price deviations to deficits in immediate cash coverage relative to outstanding supply. The coin issuer’s control mechanisms comprise two key instruments: the reallocation rate between immediately available cash and short-duration government bills, and the spread applied to mint or burn requests.
@@
The numerical experiments show that the window-based controller outperforms empirically common treasury heuristics. Relative to (i) a maximum-yield rule that defers liquidity building until the final settlement window and (ii) a maximum-liquidity rule that sacrifices carry, the controller retains a large share of bill carry in benign conditions while pre-emptively building cash as projected redemption intensities rise, thereby reducing de-peg risk. Because the policy depends only on filtered intensities, outstanding supply, and window-averaged shadow values, it fits daily treasury workflows and supports supervisory review. More broadly, the framework makes the link between balance-sheet choices and market observables explicit: peg deviations can occur only when intra-window cash is insufficient relative to the outstanding supply, and the controller’s reallocation rate and fee schedule manage that risk directly.
Several potential directions for future research emerge from the framework presented in this paper. A critical risk management concern is the tail probability of a complete depeg. Since such an event is close to catastrophic for the issuer, its probability warrants careful measurement even when it is very small. In this article we simulated 100 runs, which provides only limited estimation precision of probabilities below the one percent level. Future analyses can tighten this by running much larger simulation batches on specialized hardware, by applying dedicated rare-event techniques, or by developing analytical tail approximations.
@@
During the preparation of this manuscript, the author used ChatGPT 5 and Claude 4 for stylistic improvements and coding assistance. The outputs were carefully reviewed and revised, and all responsibility for the final content rests with the author.
@@
The author reports there are no competing interests to declare.
Funding
No funding was received for the preparation of this manuscript.